# Fingerprint Matching with Localized Deep Representation

### v1

@@ I Introduction
However, minutia-based fingerprint representation poses several challenges: (1) High computational complexity. Combining slow minutiae matching with efficient candidate filtering approaches is usually required for large-scale fingerprint identification systems[11, 12], making the entire system very complex and possibly reducing the recognition accuracy; (2) Drastic performance degradation in fingerprints of low quality. Additional fingerprint features are introduced to improve the matching performance[13, 14], while further increasing the computation and storage cost; (3) Difficult template encryption. Currently available template encryption technologies for minutia sets are far fewer than those for fixed-length representations[15], and tend to result in significant performance reduction[16]; (4) Hard to model the statistical distributions of distance between genuine or impostor pairs. This makes it difficult to obtain reliable and accurate random match probability prediction like DNA typing[17], which is very important for forensic applications[18]; (5) Difficult to implement minutiae matcher as end-to-end deep networks. Although several intermediate steps can be implemented by deep networks, such as extracting minutiae[19, 20] and their local descriptors[21], it is hard to optimize the whole fingerprint matcher if not implemented as end-to-end deep networks.
@@
Therefore, researchers have proposed several fixed-length fingerprint representations, with the aim for fingerprint indexing[22, 21, 23, 24, 25, 26, 27, 28], template security[29, 30, 31, 24], and matching low quality fingerprints[32, 33]. Conventional fixed-length fingerprint representations are usually defined by local responses of multiple Gabor filters[32], parameters of orientation filed model[34], local minutiae vicinities[35], histograms of minutia triplets or pairs features[36, 30], local ridge orientation and frequency[22], and local discretized texture descriptors[31]. The accuracies of these representations are typically much lower than those achieved by state-of-the-art (SOTA) minutiae matchers. With the rapid development of deep learning technology, more recently, fixed-length representations were extracted utilizing deep networks from aligned fingerprints[21, 24, 28]. Such methods present excellent performance on fingerprints with complete area such as rolled fingerprints. However, background area is not negligible and may dominate in fingerprints with diverse visible areas and impression types, particularly when fingerprint areas are small, resulting in a dramatic decrease in the matching performance. Although minutia information was introduced in several approaches explicitly[25, 27] and implicitly[26], the accuracy of minutia extraction itself may be compromised. Additionally, the matching is also not restricted within the overlapping area since the global aggregation step in such methods erases both the number and location information of minutiae, which limits the applicability of such methods for matching fingerprints with diverse visible areas. Extracting pyramid features at different scales and concatenating them as a feature vector was considered in[33]. However, matching performance is still affected by the coarse local region division and fusing weights for different fingerprints in such methods. Besides, extraction these multi-scale features also consumes higher computational resources and storage space.
@@ III-A2 Localized Feature Extraction
where $(x,y)$ refers to the two-dimensional coordinate indices of the feature map, $i$ is the index for channel dimension and $i\in\left[0,\left.d_{\text{model}}\middle/4\right.\right)$, and the function $t(x,i)$ or $t(y,i)$ is computed by
@@ V-B Evaluation on Fingerprint Matching
VeriFinger[57] Minutia-based
99.80
99.65
99.96
99.96
99.70
99.45
DeepPrint[24]2 Fixed-length
96.25
98.60
95.52
97.26
94.95
97.40
V-DeepPrint[24]3 Fixed-length
96.55
99.00
98.67
99.59
95.55
98.20
AFRNet[28]4 Fixed-length
94.75
97.40
97.00
97.78
96.25
96.85
V-AFRNet[28]5 Fixed-length
97.90
99.30
99.89
99.85
98.60
98.90
Multi-Scale[33] Fixed-length
97.80
98.85
99.07
99.85
97.75
98.40
Ours (binary) Fixed-length
99.35
99.40
99.85
99.85
98.55
98.50
Ours Fixed-length
99.40
99.55
99.89
99.89
99.00
98.85
Fusion -
99.80
99.75
99.96
100
@@
VeriFinger[57] Minutia-based
99.89
99.96
99.50
99.89
99.93
99.43
98.93
97.79
99.86
92.65
100
99.47
DeepPrint[24] Fixed-length
96.50
97.79
89.43
93.75
95.46
70.96
94.82
93.46
94.43
83.02
99.86
91.71
V-DeepPrint[24] Fixed-length
96.03
96.67
90.10
94.71
97.67
69.85
94.00
91.10
96.32
82.39
99.76
90.21
AFRNet[28] Fixed-length
99.18
98.82
97.14
98.04
99.46
80.25
98.14
96.04
96.46
93.31
99.99
94.57
V-AFRNet[28] Fixed-length
88.36
96.18
84.64
76.82
90.54
49.82
87.43
79.25
92.32
59.49
99.27
89.81
Multi-Scale[33] Fixed-length
99.00
99.68
91.57
98.89
99.61
83.36
97.36
94.75
94.46
47.37
99.92
92.48
Ours (binary) Fixed-length
98.92
99.29
96.11
98.61
99.61
90.75
98.57
97.39
97.54
83.89
99.97
95.15
Ours Fixed-length
98.79
99.46
97.29
99.18
99.68
94.93
98.64
98.29
97.50
88.92
100
96.09
Fusion -
99.93
100
99.50
99.93
99.93
99.46
99.32
98.07
99.86
93.41
100
99.57

### v2

@@ I Introduction
However, minutia-based fingerprint representation poses several challenges: (1) High computational complexity. Combining slow minutiae matching with efficient candidate filtering approaches is usually required for large-scale fingerprint identification systems[11, 12], making the entire system very complex and possibly reducing the recognition accuracy; (2) Drastic performance degradation in fingerprints of low quality. Additional fingerprint features are introduced to improve the matching performance[13, 14], while further increasing the computation and storage cost; (3) Difficult template encryption. Currently available template encryption technologies for minutia sets are far fewer than those for fixed-length representations[15], and tend to result in significant performance reduction[16]; (4) Hard to model the statistical distributions of distance between genuine or impostor pairs[17, 18, 19]. This makes it difficult to obtain reliable and accurate random match probability prediction like DNA typing[20], which is very important for forensic applications[21]; (5) Difficult to implement minutiae matcher as end-to-end deep networks. Although several intermediate steps can be implemented by deep networks, such as extracting minutiae[22, 23] and their local descriptors[24], it is hard to optimize the whole fingerprint matcher if not implemented as end-to-end deep networks.
@@
Therefore, researchers have proposed several fixed-length fingerprint representations, with the aim for fingerprint indexing[25, 24, 26, 27, 28, 29, 30, 31], template security[32, 33, 34, 27], and matching low quality fingerprints[35, 36]. Conventional fixed-length fingerprint representations are usually defined by local responses of multiple Gabor filters[35], parameters of orientation field model[37, 38], local minutiae vicinities[39], histograms of minutia triplets or pairs features[40, 33], local ridge orientation and frequency[25], and local discretized texture descriptors[34]. The accuracies of these representations are typically much lower than those achieved by state-of-the-art (SOTA) minutiae matchers. With the rapid development of deep learning technology, more recently, fixed-length representations were extracted utilizing deep networks from aligned fingerprints[24, 27, 31]. Such methods present excellent performance on fingerprints with complete area such as rolled fingerprints. However, background area is not negligible and may dominate in fingerprints with diverse visible areas and impression types, particularly when fingerprint areas are small, resulting in a dramatic decrease in the matching performance. Although minutia information was introduced in several approaches explicitly[28, 30] and implicitly[29], the accuracy of minutia extraction itself may be compromised. Additionally, the matching is also not restricted within the overlapping area since the global aggregation step in such methods erases both the number and location information of minutiae, which limits the applicability of such methods for matching fingerprints with diverse visible areas. Extracting pyramid features at different scales and concatenating them as a feature vector was considered in[36]. However, matching performance is still affected by the coarse local region division and fusing weights for different fingerprints in such methods. Besides, extraction these multi-scale features also consumes higher computational resources and storage space.
@@ III-A2 Localized Feature Extraction
where $(x,y)$ refers to the two-dimensional coordinate indices of the feature map, $i$ is the index for channel dimension and $i\in\left[0,\left.d_{\text{model}}\middle/4\right.\right)$,
and the function $t(x,i)$ or $t(y,i)$ is computed by
@@ V-B Evaluation on Fingerprint Matching
VeriFinger[61] Minutia-based 99.80 99.65 99.96 99.96 99.70 99.45
DeepPrint[27]2 Fixed-length 96.25 98.60 95.52 97.26 94.95 97.40
V-DeepPrint[27]3 Fixed-length 96.55 99.00 98.67 99.59 95.55 98.20
AFRNet[31]4 Fixed-length 94.75 97.40 97.00 97.78 96.25 96.85
V-AFRNet[31]5 Fixed-length 97.90 99.30 99.89 99.85 98.60 98.90
Multi-Scale[36] Fixed-length 97.80 98.85 99.07 99.85 97.75 98.40
Ours (binary) Fixed-length 99.35 99.40 99.85 99.85 98.55 98.50
Ours Fixed-length 99.40 99.55 99.89 99.89 99.00 98.85
Fusion - 99.80 99.75 99.96 100 - -
@@
VeriFinger[61] Minutia-based 99.89 99.96 99.50 99.89 99.93 99.43 98.93 97.79 99.86 92.65 100 99.47
DeepPrint[27] Fixed-length 96.50 97.79 89.43 93.75 95.46 70.96 94.82 93.46 94.43 83.02 99.86 91.71
V-DeepPrint[27] Fixed-length 96.03 96.67 90.10 94.71 97.67 69.85 94.00 91.10 96.32 82.39 99.76 90.21
AFRNet[31] Fixed-length 99.18 98.82 97.14 98.04 99.46 80.25 98.14 96.04 96.46 93.31 99.99 94.57
V-AFRNet[31] Fixed-length 88.36 96.18 84.64 76.82 90.54 49.82 87.43 79.25 92.32 59.49 99.27 89.81
Multi-Scale[36] Fixed-length 99.00 99.68 91.57 98.89 99.61 83.36 97.36 94.75 94.46 47.37 99.92 92.48
Ours (binary) Fixed-length 98.92 99.29 96.11 98.61 99.61 90.75 98.57 97.39 97.54 83.89 99.97 95.15
Ours Fixed-length 98.79 99.46 97.29 99.18 99.68 94.93 98.64 98.29 97.50 88.92 100 96.09
Fusion - 99.93 100 99.50 99.93 99.93 99.46 99.32 98.07 99.86 93.41 100 99.57